y = x + 1
2y - x = 6
The lines whose equations are shown intersect at which point?

- (0, -2s)
- (-2s, 0)
- (4s, 2s)
Please Help!(:
-Kate(:

Answer by ewatrrr(24785)

 
Hi
y = x + 1 Green
2y - x = 6 OR y =.5x+3
solving by substitution:
2(x+1) -x = 6
x + 2 = 6
x = 4 and y = 5